OpenStudy (anonymous):

simplify: m^2+m-6/2p * 8/m^2-2m i know the answer is either: A) 4(m+3)/mp, m ≠ 0, 2 and p ≠ 0 B) 4(m+3)/mp, m ≠ −2, 0 and p ≠ 0

jimthompson5910 (jim_thompson5910):

Both answers are the same in the first part. So we can ignore that portion. The second part is the only difference Since p is not equal to zero in both cases, we can ignore p and focus on m Since m^2 - 2m = 0 means that m^2 - 2m = 0 m(m - 2) = 0 m=0 or m-2=0 m=0 or m=2 These values make the denominator zero. So we must make sure that m can't equal these values So m ≠ 0 or m ≠ 2 which shortens to m ≠ 0, 2

OpenStudy (anonymous):

okay , thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome